AAVSOnet Status, 2021-07-01

Affiliation
None
Thu, 07/01/2021 - 17:39

Several changes since the last report.  Cliff is working on the photometry part of the python version of the pipeline.  George and Ken have been working diligently on creating and maintaining target plans.

BSM_Berry.  This system is operational, though there is some weirdness in the Paramount ME declination axis.  If someone has a lot of experience working with Paramounts, get in touch with Greg Bolt and give him some guidance!  Basically, pointing can be off by 10-20arcmin even though a good model is in place.  For normal operation, this doesn't have a lot of impact, as we always recenter a field after a slew, but it adds overhead.

BSM_Hamren.  There is currently a mount communication issue that we are working on.  Hopefully this will be resolved shortly.

BSM_NH2.  Fully operational.

BSM_NM.  Fully operational.  This is the Southwestern monsoonal period, so little observing will be done until September.

BSM_S.  Peter has moved, and is working on getting this system back in operation.

BSM_TX.  Fully operational.  This is the Southwestern monsoonal period, so little observing will be done until September.

OC61.  There were moisture issues with the camera.  Nigel has cleaned the sensor, and is waiting for some new dessicant packets.  We expect to be back on-line shortly.

SRO.  There were some dome shutter issues that prevented observing for the last half of June. This is the Southwestern monsoonal period, so little observing will be done until September.

TMO61.  The QHY600 camera has been installed.  However, only Sloan filters are available.  This is the Southwestern monsoonal period, so little observing will be done until September.